OpenStudy (anonymous):

If f(x)=x^2 -3x-4, find f(a+h)-f(a)

jimthompson5910 (jim_thompson5910):

Replace every x with a+h \[\Large f(x)=x^2-3x-4\] \[\Large f(\color{red}{x})=(\color{red}{x})^2-3(\color{red}{x})-4\] \[\Large f(\color{red}{a+h})=(\color{red}{a+h})^2-3(\color{red}{a+h})-4\]

jimthompson5910 (jim_thompson5910):

Now expand out and simplify \[\Large f(a+h) = (a+h)^2 -3(a+h)-4\]

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I got a^2+2ah+h^2-3a-3h-4

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

what is f(a) equal to?

OpenStudy (anonymous):

I don't know

jimthompson5910 (jim_thompson5910):

just replace every x with 'a' in f(x)

OpenStudy (anonymous):

okay thankyou

jimthompson5910 (jim_thompson5910):

\[\Large f(a+h) = a^2+2ah+h^2-3a-3h-4\] \[\Large f(a) = a^2-3a-4\]

jimthompson5910 (jim_thompson5910):

now tell me what \[\Large f(a+h)-f(a)\] is equal to

OpenStudy (anonymous):

2ah+h^2-3h

jimthompson5910 (jim_thompson5910):

correct